# Omni2Sound: Towards Unified Video-Text-to-Audio Generation

### v1

@@ 4.1 Foundation Model Architecture
We adhere to a principle of simplicity and scalability, adopting a standard Diffusion Transformer (DiT) backbone [Peebles2023DiT] conditioned on latent features from a pre-trained audio VAE [evans2024fast]. As shown in Figure 3, the backbone is conditioned on multimodal inputs using a decoupled injection approach, which is separated into two distinct branches: (1) Semantic Branch (What) and (2) Temporal Branch (When). To capture global semantic context, we concatenate text embeddings from Flan-T5 [chung2024scaling] ($F_{t}$) and visual features from CLIP [radford2021learning] ($F_{v}$, sampled at 8 fps) along the temporal dimension, which are then injected via cross-attention layers. Crucially, this design allows for flexible unimodal generation (V2A or T2A) by simply omitting the absent modality without requiring padding constraints. For the Temporal Branch, to ensure fine-grained synchronization, we follow [cheng2025mmaudio] to utilize a Synchformer [Viertola2024Synchformer] to extract dense visual-temporal features ($F_{s}$) and then inject it globally via Adaptive Layer Normalization (AdaLN).
This decoupled architecture effectively (1) achieves the flexibility of multi-condition frameworks like AudioX [cheng2025mmaudio], supporting extensible conditions without architectural modification; and (2) maintains precise temporal alignment comparable to MMAudio [cheng2025mmaudio] (powered by its well-designed MM-DiT architecture).
@@ 4.2 Three-stage Progressive Multi-task Training
We first conduct standalone T2A pretraining on large-scale text-audio pairs without a quality filter. Following the latent diffusion framework [stableaudio2024, Peebles2023DiT], our DiT backbone (Section 4.1) is trained to progressively denoise a noisy latent $z_{t}$ at timestep $t$, conditioned on text embeddings $H_{c}$. The model, $\epsilon_{\theta}$, is optimized via the simple L2 loss:
@@
Table 3: Comparison on VGGSound-Omni benchmark: Omni2Sound against SOTA models on T2A, V2A, and VT2A tasks. The w/ Video-LLaMA caps row evaluates Omni2Sound’s generalization to unseen captions generated by Video-LLaMA [zhang2023video-llama].
@@
T2A AudioX [tian2025audiox] 1.68 9.04 1.42 109.94 6.37 15.15 - - 0.49
MMAudio [cheng2025mmaudio] 1.92 8.62 1.63 101.66 5.84 14.30 - - 0.50
@@
V2A V-AURA [viertola2024vaura] 2.28 16.43 2.34 245.25 5.74 10.82 0.69 0.28 0.32
Frieren [wang2024frieren] 2.73 12.13 1.23 123.75 5.82 11.32 0.86 0.21 0.31
AudioX [tian2025audiox] 2.96 12.73 1.42 121.82 6.17 13.34 1.22 0.24 0.34
MMAudio [cheng2025mmaudio] 2.11 5.65 0.81 69.33 5.72 11.85 0.48 0.28 0.43
@@
VT2A ThinkSound ($w/o$. CoT) [liu2025thinksound] 1.60 7.41 1.10 116.08 6.21 11.73 0.53 0.26 0.43
HunyuanVideo-Foley [chen2024hunyuanfoley] 1.74 10.02 2.36 100.53 6.18 11.58 0.57 0.32 0.45
AudioX [tian2025audiox] 1.59 8.29 1.24 103.37 6.17 14.94 1.23 0.26 0.49
MMAudio [cheng2025mmaudio] 1.63 5.28 0.91 68.44 5.84 13.44 0.49 0.29 0.49
@@ 5 VGGSound-Omni: Unified Evaluation
A significant challenge in evaluating unified Video-Text-to-Audio (VT2A) models is the absence of a comprehensive benchmark. The VGGSound test set [chen2020vggsound] only provides sparse event labels and lacks detailed captions. Although recent work like VGGSounder [zverev2025vggsounder] significantly improved this by correcting and adding crucial modality labels (e.g., A, V, AV) for fidelity evaluation, it still lacks human-expert-level captions. To address this gap, we construct VGGSound-Omni, a new multi-track benchmark derived from the original VGGSound test set, designed for both standard unified and specialized off-screen VT2A tasks evaluations. The construction process is detailed below.
@@
Our first step was to establish a high-fidelity, human-level caption set for all 14,000 videos, forming the primary evaluation track. We first generated an initial caption using our agentic pipeline (Section 3). We then systematically validated this output via an AI-assisted verification workflow: GPT-5 [openai2025gpt5intro] was tasked to act as an auditor, checking if our captions semantically covered all the “A” and “AV” labels from VGGSounder [zverev2025vggsounder]. Samples flagged with a mismatch were routed for targeted human verification. During this manual audit process, we found most of these flagged discrepancies stemmed from annotation errors within the VGGSounder data itself (e.g., label redundancy and human annotation errors caused by visual interference). After manually correcting for these identified errors, we established our final, human-verified captions as the definitive ground truth (GT) for evaluating all three tasks (VT2A, V2A, and T2A).
Complementing the primary set, we construct a challenging off-screen track (1,048 items). We curated this subset from two sources: (i) Natural events, filtered from VGGSound for low A-V correspondence (via IB-Score [girdhar2023imagebind] and Desync-Score [cheng2025mmaudio]) while excluding background speech; and (ii) Synthetic music, formed by mixing aligned background clips from MusicCaps [agostinelli2023musiccaps]. More Details are provided in Appendix D.
@@
T2A AudioX [tian2025audiox] 2.73 19.43 3.32 171.60 5.98 12.15 - - 0.28
MMAudio [cheng2025mmaudio] 2.54 11.25 5.07 142.71 5.54 9.28 - - 0.28
@@
V2A AudioX [tian2025audiox] 3.13 18.90 4.01 205.48 5.87 8.31 1.20 0.23 0.13
MMAudio [cheng2025mmaudio] 2.94 13.41 3.87 159.30 5.50 7.59 0.62 0.24 0.14
@@
VT2A ThinkSound ($w/o$. CoT) [liu2025thinksound] 2.53 11.99 3.52 206.93 5.77 6.09 0.66 0.22 0.19
HunyuanVideo-Foley [chen2024hunyuanfoley] 2.13 8.06 3.58 94.64 6.04 8.17 0.55 0.34 0.23
AudioX [tian2025audiox] 2.39 14.26 3.16 149.37 5.97 10.23 1.21 0.23 0.26
MMAudio [cheng2025mmaudio] 2.41 10.12 4.90 129.21 5.53 7.46 0.59 0.25 0.20
@@ 6.1 Experiment Settings
For T2A backbone pre-training, we use a large-scale corpus comprising the train set of audio datasets such as AudioCaps [kim2019audiocaps], WavCaps [mei2023wavcaps], Clotho [drossos2020clotho], AudioSet [gemmeke2017audio], VGGSound [chen2020vggsound], FSD50k [fonseca2022fsd50k], as well as music datasets including MSD [bertin2011million] and FMA [defferrard2016fma]. To maintain consistency, all audio is segmented into 10-second clips and resampled at 16 kHz. Following this, the model is trained for unified VT2A tasks using our proposed SoundAtlas (Section 5) and a high-quality, PQ-score-filtered T-A subset derived from the aforementioned pre-training corpus. More details of the implementation are provided in Appendix Section G. For evaluation, we compare Omni2Sound with SOTA models on three benchmarks: our proposed VGGSound-Omni (Section 5), Kling-Audio-Eval [wang2025audiogenomni] and AudioCaps test set [kim2019audiocaps]. We strictly ensure that these evaluation benchmarks are strictly disjoint from all data used in our training stages to prevent potential data leakage.
@@
We implement our objective evaluation using the standardized AV-benchmark toolkit [hkchengrex2024avbenchmark] on 8-second clips, following previous work [cheng2025mmaudio]. We assess quality across four critical dimensions [liu2023audioldm]. For Distribution Matching, we measure feature similarity between generated and ground-truth audio using Fréchet Distance (FAD [hershey2017cnn], $\mathrm{FD}_{\mathrm{PaSST}}$ [koutini2022passt], FD [kong2020panns]) and Kullback-Leibler divergence (KL, $\mathrm{KL}_{\mathrm{PaSST}}$). Audio Quality is assessed via Inception Scores (IS [salimans2016improved], $\mathrm{IS}_{\mathrm{PaSST}}$) and Production Quality (PQ [tjandra2025meta]) for aesthetics. Semantic Alignment evaluates text-audio consistency (CLAP [elizalde2023clap], MS-CLAP [wu2023large]) and video-audio alignment (IB [girdhar2023imagebind]). Finally, Temporal Alignment is measured using the Desynchronization Score (DS) predicted by Synchformer [iashin2024synchformer]. Detailed metric definitions and calculations are provided in the Appendix.
@@ 6.2 Main Results
We present our main results on VGGSound-Omni benchmark in Table 3. To ensure a fair comparison, all baseline models are re-evaluated using their official checkpoints and the standardized AV-benchmark toolkit [hkchengrex2024avbenchmark], using the same video and text conditions. The results demonstrate that Omni2Sound achieves state-of-the-art performance across all three unified tasks (T2A, V2A, and VT2A) compared to both previous unified VT2A models (AudioX [tian2025audiox], MMAudio [cheng2025mmaudio]) and specialized models (e.g. ThinkSound [liu2025thinksound], HunyuanVideo-Foley [chen2024hunyuanfoley]). To further validate Omni2Sound’s generalization beyond our SoundAtlas captioning style, we evaluate it on the same VGGSound test clips but use the Video-LLaMA [zhang2023video-llama] captions from ThinkSound [liu2025thinksound]. As shown in Table 2 (w/ Video-LLaMA caps), while performance sees a slight degradation, our model’s scores still surpass all baselines, confirming its robustness to unseen captioning styles.
@@
To validate generalization, we evaluate on Kling-Audio-Eval [wang2025audiogenomni] and AudioCaps [kim2019audiocaps] results in Table 4 and Appendix Table 7. On Kling-Audio-Eval, Omni2Sound remains highly competitive despite the domain gap (YouTube-sourced SoundAtlas vs. Kling’s professional video). While trailing HunyuanVideo-Foley [chen2024hunyuanfoley] in some metrics, which is expected given its massive data advantage (100k vs 2k hours), our model consistently outperforms other unified and specialized baselines across all tasks. Furthermore, on AudioCaps, Omni2Sound achieves top-tier performance against specialized T2A models, securing the best scores in distribution metrics ($\mathrm{KL}$, $\mathrm{FD}$) and semantic alignment ($\mathrm{CLAP}=0.36$), while remaining highly competitive in audio quality (PQ) and the $\mathrm{FAD}$ metric.
@@ Appendix A Cost Analysis on Audio Captioning
While Gemini 2.5 Pro [team2023gemini] represents a milestone as a native multimodal foundation model, utilizing it directly for large-scale video-grounded audio captioning proves economically unsustainable. As quantified in Table 6, using Gemini’s standard API pricing, a naive implementation—processing raw video frames alongside audio (V+AV+A)—incurs a prohibitive expenditure of \$10,275 USD per 1M samples. This figure is derived from the token consumption of a 10-second sample: the input aggregates to 3,820 tokens (comprising 1,000 instruction, 320 audio, and 2,500 visual tokens), while the full chain-of-thought generation requires $\sim$550 output tokens. Crucially, this naive approach suffers from an inherent visual bias, as shown in Figure 1 in main paper.
@@
AudioLDM 2-L [liu2024audioldm2] 1.73 34.21 2.26 5.93 0.24
TANGO 2 [majumder2024tango2] 1.19 15.92 3.17 5.82 0.35
Make-An-Audio 2 [huang2023makeanaudio2_placeholder] 1.38 15.34 1.46 5.64 0.25
GenAU-Large [haji-ali2024genau] 1.42 16.92 1.32 5.52 0.26
MMAudio [cheng2025mmaudio] 1.43 13.78 2.92 5.30 0.29
AudioX [tian2025audiox] 1.55 17.10 2.65 5.81 0.31
@@ Appendix D Off-Screen Track of VGGSound-Omni
To address the high demand for Background Music (BGM) generation, we create a Synthetic Music subset by mixing semantically aligned MusicCaps [agostinelli2023musiccaps] clips into a pool of high-fidelity videos. This process follows a two-stage procedure. In the Base Selection stage, we first select a "clean" video pool by strictly requiring a 100% AV label ratio and filtering for high alignment (ImageBind $\geq 0.30$, Desync <0.55<0.55), ensuring all original acoustic events are visually manifest. Subsequently, during Semantic Mixing, we augment these videos with background music tracks. To guarantee semantic coherence, we utilize GPT to retrieve the most congruent music track from a random candidate batch of 50 samples based on the video context. Ground-truth captions are updated to reflect this acoustic addition.
@@
We acknowledge the pioneering work of VinTAGe-Bench [bain2025vintage] in synthetic robustness evaluation. However, the off-screen subset of our VGGSound-Omni benchmark extends this direction in three critical dimensions. First, in terms of Realism, by leveraging VGGSounder [zverev2025vggsounder] metadata, our natural subset is primarily sourced from real-world off-screen audio events rather than relying solely on synthetic mixes. Second, regarding Scale, our benchmark is significantly larger, providing 1,613 evaluation items compared to the 212 basic videos of VinTAGe-Bench. Third, regarding Scope, we include a dedicated Synthetic Music (BGM) track, addressing a critical, high-demand scenario often overlooked in standard environmental sound benchmarks.
@@
To further validate our model’s generalization and mitigate potential biases from our self-constructed benchmark, we evaluate it on the Kling-Audio-Eval [wang2025audiogenomni] and Audiocaps test set [kim2019audiocaps]. In Table 4, on the Kling-Audio-Eval benchmark, Omni2Sound remains highly competitive, despite a significant data scale and distribution gap (our YouTube-sourced SoundAtlas vs. Kling’s professional video/Foley). While HunyuanVideo-Foley [chen2024hunyuanfoley] leads on several metrics, this is expected given its massive 100k-hour internal dataset, which is tens of times larger than our SoundAtlas filter derived from VGGSound and AudioSet. Nevertheless, Omni2Sound consistently outperforms all other strong baselines (e.g., MMAudio, AudioX, and ThinkSound) across V2A and VT2A tasks, demonstrating strong generalization as the SOTA or second-best method. In Table 7, on the Audiocaps test set, we compare Omni2Sound against specialized SOTA T2A models. The results show our unified model achieves top-tier performance, attaining the best scores in key distribution metrics ($\mathrm{KL}$, $\mathrm{FD}$) and semantic alignment ($\mathrm{CLAP}=0.36$), while remaining highly competitive in audio quality (PQ) and the $\mathrm{FAD}$ metric.

### v2

@@ 4.1 Foundation Model Architecture
We adhere to a principle of simplicity and scalability, adopting a standard Diffusion Transformer (DiT) backbone [37] conditioned on latent features from a pre-trained audio VAE [12]. As shown in Figure 3, the backbone is conditioned on multimodal inputs using a decoupled injection approach, which is separated into two distinct branches: (1) Semantic Branch (What) and (2) Temporal Branch (When). To capture global semantic context, we concatenate text embeddings from Flan-T5 [8] ($F_{t}$) and visual features from CLIP [38] ($F_{v}$, sampled at 8 fps) along the temporal dimension, which are then injected via cross-attention layers. Crucially, this design allows for flexible unimodal generation (V2A or T2A) by simply omitting the absent modality without requiring padding constraints. For the Temporal Branch, to ensure fine-grained synchronization, we follow [7] to utilize a Synchformer [23] to extract dense visual-temporal features ($F_{s}$) and then inject it globally via Adaptive Layer Normalization (AdaLN).
This decoupled architecture effectively (1) achieves the flexibility of multi-condition frameworks like AudioX [7], supporting extensible conditions without architectural modification; and (2) maintains precise temporal alignment comparable to MMAudio [7] (powered by its well-designed MM-DiT architecture).
@@ 4.2 Three-stage Progressive Multi-task Training
We first conduct standalone T2A pretraining on large-scale text-audio pairs without a quality filter. Following the latent diffusion framework [13, 37], our DiT backbone (Section 4.1) is trained to progressively denoise a noisy latent $z_{t}$ at timestep $t$, conditioned on text embeddings $H_{c}$. The model, $\epsilon_{\theta}$, is optimized via the simple L2 loss:
@@
Table 3: Comparison on VGGSound-Omni benchmark: Omni2Sound against SOTA models on T2A, V2A, and VT2A tasks. The w/ Video-LLaMA caps row evaluates Omni2Sound’s generalization to unseen captions generated by Video-LLaMA [52].
@@
T2A AudioX [43] 1.68 9.04 1.42 109.94 6.37 15.15 - - 0.49
MMAudio [7] 1.92 8.62 1.63 101.66 5.84 14.30 - - 0.50
@@
V2A V-AURA [45] 2.28 16.43 2.34 245.25 5.74 10.82 0.69 0.28 0.32
Frieren [47] 2.73 12.13 1.23 123.75 5.82 11.32 0.86 0.21 0.31
AudioX [43] 2.96 12.73 1.42 121.82 6.17 13.34 1.22 0.24 0.34
MMAudio [7] 2.11 5.65 0.81 69.33 5.72 11.85 0.48 0.28 0.43
@@
VT2A ThinkSound ($w/o$. CoT) [32] 1.60 7.41 1.10 116.08 6.21 11.73 0.53 0.26 0.43
HunyuanVideo-Foley [40] 1.74 10.02 2.36 100.53 6.18 11.58 0.57 0.32 0.45
AudioX [43] 1.59 8.29 1.24 103.37 6.17 14.94 1.23 0.26 0.49
MMAudio [7] 1.63 5.28 0.91 68.44 5.84 13.44 0.49 0.29 0.49
@@ 5 VGGSound-Omni: Unified Evaluation
A significant challenge in evaluating unified Video-Text-to-Audio (VT2A) models is the absence of a comprehensive benchmark. The VGGSound test set [4] only provides sparse event labels and lacks detailed captions. Although recent work like VGGSounder [54] significantly improved this by correcting and adding crucial modality labels (e.g., A, V, AV) for fidelity evaluation, it still lacks human-expert-level captions. To address this gap, we construct VGGSound-Omni, a new multi-track benchmark derived from the original VGGSound test set, designed for both standard unified and specialized off-screen VT2A tasks evaluations. The construction process is detailed below.
@@
Our first step was to establish a high-fidelity, human-level caption set for all 14,000 videos, forming the primary evaluation track. We first generated an initial caption using our agentic pipeline (Section 3). We then systematically validated this output via an AI-assisted verification workflow: GPT-5 [openai2025gpt5intro] was tasked to act as an auditor, checking if our captions semantically covered all the “A” and “AV” labels from VGGSounder [54]. Samples flagged with a mismatch were routed for targeted human verification. During this manual audit process, we found most of these flagged discrepancies stemmed from annotation errors within the VGGSounder data itself (e.g., label redundancy and human annotation errors caused by visual interference). After manually correcting for these identified errors, we established our final, human-verified captions as the definitive ground truth (GT) for evaluating all three tasks (VT2A, V2A, and T2A).
Complementing the primary set, we construct a challenging off-screen track (1,048 items). We curated this subset from two sources: (i) Natural events, filtered from VGGSound for low A-V correspondence (via IB-Score [17] and Desync-Score [7]) while excluding background speech; and (ii) Synthetic music, formed by mixing aligned background clips from MusicCaps [1]. More Details are provided in Appendix D.
@@
T2A AudioX [43] 2.73 19.43 3.32 171.60 5.98 12.15 - - 0.28
MMAudio [7] 2.54 11.25 5.07 142.71 5.54 9.28 - - 0.28
@@
V2A AudioX [43] 3.13 18.90 4.01 205.48 5.87 8.31 1.20 0.23 0.13
MMAudio [7] 2.94 13.41 3.87 159.30 5.50 7.59 0.62 0.24 0.14
@@
VT2A ThinkSound ($w/o$. CoT) [32] 2.53 11.99 3.52 206.93 5.77 6.09 0.66 0.22 0.19
HunyuanVideo-Foley [40] 2.13 8.06 3.58 94.64 6.04 8.17 0.55 0.34 0.23
AudioX [43] 2.39 14.26 3.16 149.37 5.97 10.23 1.21 0.23 0.26
MMAudio [7] 2.41 10.12 4.90 129.21 5.53 7.46 0.59 0.25 0.20
@@ 6.1 Experiment Settings
For T2A backbone pre-training, we use a large-scale corpus comprising the train set of audio datasets such as AudioCaps [24], WavCaps [36], Clotho [10], AudioSet [15], VGGSound [4], FSD50k [14], as well as music datasets including MSD [3] and FMA [9]. To maintain consistency, all audio is segmented into 10-second clips and resampled at 16 kHz. Following this, the model is trained for unified VT2A tasks using our proposed SoundAtlas (Section 5) and a high-quality, PQ-score-filtered T-A subset derived from the aforementioned pre-training corpus. More details of the implementation are provided in Appendix Section G. For evaluation, we compare Omni2Sound with SOTA models on three benchmarks: our proposed VGGSound-Omni (Section 5), Kling-Audio-Eval [46] and AudioCaps test set [24]. We strictly ensure that these evaluation benchmarks are strictly disjoint from all data used in our training stages to prevent potential data leakage.
@@
We implement our objective evaluation using the standardized AV-benchmark toolkit [7] on 8-second clips, following previous work [7]. We assess quality across four critical dimensions [30]. For Distribution Matching, we measure feature similarity between generated and ground-truth audio using Fréchet Distance (FAD [19], $\mathrm{FD}_{\mathrm{PaSST}}$ [26], FD [25]) and Kullback-Leibler divergence (KL, $\mathrm{KL}_{\mathrm{PaSST}}$). Audio Quality is assessed via Inception Scores (IS [39], $\mathrm{IS}_{\mathrm{PaSST}}$) and Production Quality (PQ [44]) for aesthetics. Semantic Alignment evaluates text-audio consistency (CLAP [11], MS-CLAP [48]) and video-audio alignment (IB [17]). Finally, Temporal Alignment is measured using the Desynchronization Score (DS) predicted by Synchformer [22]. Detailed metric definitions and calculations are provided in the Appendix.
@@ 6.2 Main Results
We present our main results on VGGSound-Omni benchmark in Table 3. To ensure a fair comparison, all baseline models are re-evaluated using their official checkpoints and the standardized AV-benchmark toolkit [7], using the same video and text conditions. The results demonstrate that Omni2Sound achieves state-of-the-art performance across all three unified tasks (T2A, V2A, and VT2A) compared to both previous unified VT2A models (AudioX [43], MMAudio [7]) and specialized models (e.g. ThinkSound [32], HunyuanVideo-Foley [40]). To further validate Omni2Sound’s generalization beyond our SoundAtlas captioning style, we evaluate it on the same VGGSound test clips but use the Video-LLaMA [52] captions from ThinkSound [32]. As shown in Table 2 (w/ Video-LLaMA caps), while performance sees a slight degradation, our model’s scores still surpass all baselines, confirming its robustness to unseen captioning styles.
@@
To validate generalization, we evaluate on Kling-Audio-Eval [46] and AudioCaps [24] results in Table 4 and Appendix Table 7. On Kling-Audio-Eval, Omni2Sound remains highly competitive despite the domain gap (YouTube-sourced SoundAtlas vs. Kling’s professional video). While trailing HunyuanVideo-Foley [40] in some metrics, which is expected given its massive data advantage (100k vs 2k hours), our model consistently outperforms other unified and specialized baselines across all tasks. Furthermore, on AudioCaps, Omni2Sound achieves top-tier performance against specialized T2A models, securing the best scores in distribution metrics ($\mathrm{KL}$, $\mathrm{FD}$) and semantic alignment ($\mathrm{CLAP}=0.36$), while remaining highly competitive in audio quality (PQ) and the $\mathrm{FAD}$ metric.
@@ Appendix A Cost Analysis on Audio Captioning
While Gemini 2.5 Pro [42] represents a milestone as a native multimodal foundation model, utilizing it directly for large-scale video-grounded audio captioning proves economically unsustainable. As quantified in Table 6, using Gemini’s standard API pricing, a naive implementation—processing raw video frames alongside audio (V+AV+A)—incurs a prohibitive expenditure of \$10,275 USD per 1M samples. This figure is derived from the token consumption of a 10-second sample: the input aggregates to 3,820 tokens (comprising 1,000 instruction, 320 audio, and 2,500 visual tokens), while the full chain-of-thought generation requires $\sim$550 output tokens. Crucially, this naive approach suffers from an inherent visual bias, as shown in Figure 1 in main paper.
@@
AudioLDM 2-L [31] 1.73 34.21 2.26 5.93 0.24
TANGO 2 [35] 1.19 15.92 3.17 5.82 0.35
Make-An-Audio 2 [20] 1.38 15.34 1.46 5.64 0.25
GenAU-Large [18] 1.42 16.92 1.32 5.52 0.26
MMAudio [7] 1.43 13.78 2.92 5.30 0.29
AudioX [43] 1.55 17.10 2.65 5.81 0.31
@@ Appendix D Off-Screen Track of VGGSound-Omni
To address the high demand for Background Music (BGM) generation, we create a Synthetic Music subset by mixing semantically aligned MusicCaps [1] clips into a pool of high-fidelity videos. This process follows a two-stage procedure. In the Base Selection stage, we first select a "clean" video pool by strictly requiring a 100% AV label ratio and filtering for high alignment (ImageBind $\geq 0.30$, Desync <0.55<0.55), ensuring all original acoustic events are visually manifest. Subsequently, during Semantic Mixing, we augment these videos with background music tracks. To guarantee semantic coherence, we utilize GPT to retrieve the most congruent music track from a random candidate batch of 50 samples based on the video context. Ground-truth captions are updated to reflect this acoustic addition.
@@
We acknowledge the pioneering work of VinTAGe-Bench [29] in synthetic robustness evaluation. However, the off-screen subset of our VGGSound-Omni benchmark extends this direction in three critical dimensions. First, in terms of Realism, by leveraging VGGSounder [54] metadata, our natural subset is primarily sourced from real-world off-screen audio events rather than relying solely on synthetic mixes. Second, regarding Scale, our benchmark is significantly larger, providing 1,613 evaluation items compared to the 212 basic videos of VinTAGe-Bench. Third, regarding Scope, we include a dedicated Synthetic Music (BGM) track, addressing a critical, high-demand scenario often overlooked in standard environmental sound benchmarks.
@@
To further validate our model’s generalization and mitigate potential biases from our self-constructed benchmark, we evaluate it on the Kling-Audio-Eval [46] and Audiocaps test set [24]. In Table 4, on the Kling-Audio-Eval benchmark, Omni2Sound remains highly competitive, despite a significant data scale and distribution gap (our YouTube-sourced SoundAtlas vs. Kling’s professional video/Foley). While HunyuanVideo-Foley [40] leads on several metrics, this is expected given its massive 100k-hour internal dataset, which is tens of times larger than our SoundAtlas filter derived from VGGSound and AudioSet. Nevertheless, Omni2Sound consistently outperforms all other strong baselines (e.g., MMAudio, AudioX, and ThinkSound) across V2A and VT2A tasks, demonstrating strong generalization as the SOTA or second-best method. In Table 7, on the Audiocaps test set, we compare Omni2Sound against specialized SOTA T2A models. The results show our unified model achieves top-tier performance, attaining the best scores in key distribution metrics ($\mathrm{KL}$, $\mathrm{FD}$) and semantic alignment ($\mathrm{CLAP}=0.36$), while remaining highly competitive in audio quality (PQ) and the $\mathrm{FAD}$ metric.